OpenStudy (anonymous):

A bag contains four red marbles, eight green marbles, three blue marbles, and nine yellow marbles. If Lisa draws a random marble from the bag, what is the probability that it will be a red, green, or blue marble?

OpenStudy (anonymous):

red = 4/20 2/10 1/5 <--- i reduced them green= 8/16 4/8 2/4 1/2 <---- blue = 3/21 1/7 <----

OpenStudy (anonymous):

1/8?

OpenStudy (anonymous):

I mean for like the total? alll of them together probability

OpenStudy (anonymous):

4/24 2/12 1/8 so yeah 1/8 is correct!

OpenStudy (anonymous):

can I ask u another?

OpenStudy (anonymous):

sure why not

OpenStudy (anonymous):

Holly is taking a multiple choice reading test. She has decided to guess on questions 3 and 7. On each question, there are four worded choices plus an "all of these" choice and a "none of these" choice. What is the probability that she answers question 3 correctly and question 7 incorrectly?

OpenStudy (anonymous):

When there are 6 choices, and 5 of them are INcorrect P(INcorrect) on each question if 5/6. To get the overall probability for two questions you multiply the individual probabilities. 5/6 * 5/6 = 25/36 chance of getting both wrong. To check you can compute the remaining cases: Both right: 1/6 * 1/6 = 1/36 first right, second wrong: 1/6 * 5/6 = 5/36 same for other way around = 5/36 Total = 36/36 = 1 as it should be.

OpenStudy (anonymous):

thanks! I was thinking it would be 5/36

OpenStudy (anonymous):

u are welcome!